OpenStudy (anonymous):

Is the following always, sometimes or never true? x + 1 = x + 0

OpenStudy (anonymous):

1=0, never true

OpenStudy (anonymous):

\[ x + 1 = x + 0 \\ x -x +1 = x -x+ 0\\1=0 \]never